What is it a logical theory of truth?

Serban Leoca

Truth is a semantic concept. A theory of truth should be a semantic theory. What special features does it have among other semantic theories? We take into consideration the truth theories of Tarski, Kripke, Barwise and Etchemendy, and Davidson’s project, and get the following answer:

A semantic theory becomes a logical theory of truth if it provides a construction of the concept of a true sentence (i.e. an extensional definition of the concept of truth) that explicitly gives an account of self-reference and, thereby, gives a solution to avoid the liar paradox.

Logical theories of truth have emerged as an answer to the challenge of the natural language: their attempt was either to elude the natural language by proposing artificial languages instead (Carnap, Tarski), or to refine artificial languages with a view to formalizing ever increasing parts of the natural language (Davidson, Kripke).

Address: Faculty of Philosophy, University of Bucharest, Romania

E-mail: sleoca@ccd.ubbcluj.ro